# Robust Containment Queries over Collections of Trimmed NURBS Surfaces via Generalized Winding Numbers

### v2

@@ 4.1. Numerical Integration of Boundary Formulation
We show the results of this comparison in the “Non-Adaptive Quadrature” row of Figure 9, where errors in the 1D scheme are only present near the surface boundary, while errors in the 2D scheme are present for all points near the surface. The accuracy of this 2D scheme, and in some sense even its applicability, would be further compromised by the presence of trimming curves, which are far more naturally handled in our boundary scheme. While it is true that both types of errors can be mitigated through the use of adaptive quadrature, applying such a strategy to a 2D scheme would require unreasonably many surface evaluations to achieve the same level of accuracy as our adaptive 1D scheme (as can be seen in the “Adaptive Quadrature” row of Figure 9).
It is this type of surface-based subdivision strategy which is described as “adaptive quadrature” and compared to in Martens and Bessmeltsev (2025), which largely explains their characterization that quadrature is inapplicable for complex CAD surfaces. In contrast, as can be seen in Figure 10, our adaptive quadrature strategy outperforms the reported metrics of that work on similar shapes, and is capable of doing so even without taking advantage of per-query parallelism, the planar characteristics of the input boundary curves, or any structure in the input set of query points. Figure 10 also demonstrates on practical examples how algorithm performance is influenced by the position of the query point. As expected, “near” points close to surface interiors have negligible overhead compared to “far” points, while only those close to a surface boundary are more expensive since they must often be treated as edge-cases.
@@ 4.2. Accuracy Evaluation
We now consider the full implementation of our algorithm, which solves Equation 7 at arbitrary points up to arbitrary accuracy with a more sophisticated, error-controlled adaptive quadrature scheme. In Figure 11, we demonstrate the versatility of our algorithm by evaluating the GWN on several models adapted from the ABC dataset (Koch et al., 2019), a collection of CAD models of varying quality used to benchmark algorithms in computer graphics and machine learning. In order to exaggerate the effects of the GWN field on open 3D shapes, we have intentionally made each of these shapes non-watertight through the deletion of several surfaces. We present qualitative results on additional watertight and non-watertight CAD models in the supplemental materials to this manuscript.
Far from open boundaries, the GWN field neatly partitions the domain into exterior and interior regions. Closer to these gaps in the shape, however, the GWN field degrades to noninteger values, corresponding to increasing uncertainty in the subsequent containment classification. The simplest strategy maps these fractional values to a containment decision by rounding them to the nearest integer and applying either a non-zero or even-odd rule. The resulting interior region is consistent with the provided boundary and provides a reasonable classification for points which are near open edges. We can see this directly in the right-most column of Figure 11, which shows how rounding the scalar GWN field “corrects” small errors in the surface by defining an implicit boundary through the $0.5$-isosurface.
@@
In Figure 12 we compare our proposed algorithm to the above adaptations of conventional alternatives on subsets of a watertight “Gear” shape, with each view considered on a structured grid of $\times500$ sample points. In Figure 12(top), we consider the shape when it is coarsely, yet still meaningfully, discretized according to each method to see the primary characteristics of the resulting GWN field. For the triangulation method, we see, as expected, that curved surfaces are made linear, resulting in unpredictable errors near the surface boundary while still producing integer-valued winding numbers for this watertight model. In contrast, while the 0.5-isosurface of scalar field provided by the point cloud method generally captures the curved features, the GWN field in its totality diverges significantly from the integer-valued ground truth. Although the dipoles correctly place nearby query points on the correct side of the curved surface, their large magnitude results in scalar GWN values outside the range of [0.5,1.5][-0.5,1.5]. This would lead to further misclassifications if containment were determined with the conventional strategy of rounding, followed by an even-odd or non-zero rule, which might be necessary on more complex shapes with nested surfaces.
@@
To see this error more quantitatively, we record in Figure 12(bottom) the total number of misclassifications on increasingly narrow views of the shape using the triangulation approach (left) and the point cloud approach (right), with each misclassified point in each $\times500$ grid being marked in red. While the number of misclassifications for each discretization strategy (generally) decreases as the resolution of the discretization increases, there is no level at which no misclassifications occur. Further, although the linear deflection parameter controls how far the triangulation can differ from the original shape, in effect ensuring that no points within a fixed absolute distance of the surface will be misclassified, it can be difficult to know a priori what level of discretization is necessary for correct classification at user-specified query points. This is well-illustrated in the narrowest view, in which the increased deflection parameter from $10^{-3}$ m to $10^{-5}$ m does not necessarily increase the accuracy of the method, as all points which are misclassified are by coincidence still closer than $10^{-5}$ m to the surface. This relationship is further obfuscated in the case of a point cloud representation, where it is similarly difficult to gauge the appropriate sampling density for a point cloud that will lead to accurate classifications. In either case, this can lead to unnecessary computational expense in the form of globally dense triangulations or sampling procedures. In contrast, operating directly on the trimmed surface allows our method to perform adaptively for each query to the requested level of accuracy and correctly classify the containment of all query points.
@@
Shape Bobbin Trailer Pipe Spring
Spring,
Two-Patch
@@ 4.3. Performance Evaluation
Other important factors come from more general characteristics of the component NURBS patches are best illustrated through comparison of the “Trailer” and “Pipe” examples. Although the Trailer shape has more surfaces, the NURBS patches themselves are predominantly flat with simple trimming curves. In contrast, the Pipe shape is more complicated, with many of its component NURBS patches being large spheres and cylinders which are trimmed and filleted. Thus, while the times to process the individual far- and near-field cases are comparable, resolving the edge-cases of the Pipe example is more challenging, and therefore more expensive. We present similar timing tests on other watertight CAD shapes from the ABC dataset in the supplemental materials to this manuscript.
@@ 5.1.1. Line-Surface Intersections
We use a geometric approach to identify the isoparametric coordinates of all intersections between a line and an untrimmed NURBS patch in the parameter space of the patch. This involves recursively subdividing the patch geometry (either through Bézier extraction of NURBS patches or bisection of individual Bézier patches) and clipping away portions of the surface where intersections provably cannot occur. By maintaining a record of each subdivision patch’s placement within the parameter space of the original surface, this method converges linearly to the set of parametric points $(u,v)$ at which the line intersects the untrimmed NURBS patch.
@@ 5.1.2. Adaptive Quadrature and Disk Radius
As implemented in our algorithm, memoized quadrature nodes cannot be reused if a patch’s trimming curves change during evaluation of the GWN. As a result, there is a strong correlation between overall computational performance and the choice of the disk radius used to identify and resolve near-field and edge-case configurations. A smaller disk radius leads to fewer points getting resolved as edge-cases, as it is less likely for the disk to intersect with existing trimming curves, which better utilizes our cached quadrature nodes. On the other hand, a smaller disk radius allows query points to be closer to the singularity in the boundary curve integrals, necessitating a stricter tolerance for $\epsilon_{q}$ to ensure consistent accuracy.
To demonstrate how both the disk radius (measured as a percent of each patch’s parameter space bounding box diagonal) and $\epsilon_{q}$ affects the accuracy and computational efficiency of the algorithm, we repeat the numerical experiment for $\epsilon_{ls}$ in Section 5.1.1 and report the results in Figure 14, varying $\epsilon_{q}$ during the evaluation of the GWN for disks of various size. In these examples, we set $\epsilon_{ls}=10^{-6}$, and further note that $\epsilon_{ls}$ and the disk radius have an independent effect on the method’s accuracy. We see that as the radius decreases, a more strict tolerance for $\epsilon_{q}$ is needed to ensure there are exceptionally few misclassifications. Furthermore, we see that across disk radii, a stricter tolerance involves a greater computational cost, but this cost is largely marginal relative to the overall effect of disk radius. It is for this reason that we use for our algorithm a disk radius that is 1% of the parameter space bounding box diagonal, as it is small enough to allow full utilization of the memoization scheme, while still allowing for a somewhat lenient tolerance of $\epsilon_{q}=10^{-6}$.
@@ 5.2. Robustness to Defective Trimming Curves
First is robustness to errors between the (valid) trimmed NURBS surfaces in the model, which is necessary when numerical gaps and overlaps between adjacent patches make conventional containment techniques like ray-casting inaccurate. By the linearity of the GWN, our algorithm is evaluated independently for each patch, ensuring our containment decision is reasonable for this kind of messy input. We achieve this form of robustness through accurate and precise calculation of the GWN field at arbitrary 3D points, one surface at a time. This property of our algorithm is exhibited on watertight and non-watertight examples throughout this paper (e.g. Figures 11011 and 12) as well as in the supplement.
@@
Figure 16 demonstrates how these decisions impact our calculation of the GWN field in the presence of a surface with poorly defined trimming curves, and in turn how this GWN field impacts the subsequent containment decision. To the original shape in Figure 16(top left), we perturb the trimming curves of the front face, but do so such that adjacent curves remain connected. The resulting set of trimming curves, shown in Figure 16(bottom left), is still closed and connected, but the perturbation has introduced self-intersections that form small clockwise loops. The boundary formulation itself is indifferent to trimming curve orientation, and so for query points away from these self-intersections, Algorithm 1 continues to produce a GWN that degrades smoothly around the mismatch with the new boundary. While this type of trimming curve geometry poses a challenge for our correction term (Equation 7) and edge-case strategy, resulting in a noisier GWN field, the containment decisions remain consistent with those of the original model.
Next, we consider collections of non-watertight trimming curves. For Figure 16(top right) we delete trimming curves at each corner in parameter space, and for Figure 16(bottom right) we delete a middle portion from each curve. In the presence of these defective trimming curves, we observe that the computed GWN field does not degrade smoothly in the same way that it does around open, yet well-defined edges. For example, applying our boundary formulation to the same malformed surface returns slightly different GWN values at the same query point depending on the chosen form of the line integral. This is seen most prominently in the boundary between regions handled by different cases, such between far-field and edge-case scenarios. Nevertheless, we see that these discontinuities are concentrated near geometric errors, such that rounding and applying an even-odd rule to the calculated GWN field still achieves excellent containment decisions for these examples.
@@ 6. Conclusions
In this work, we have introduced the first algorithm to accurately and precisely compute generalized winding numbers over unstructured collections of trimmed NURBS patches. We have used this to define a robust containment query that can be applied to general CAD models, even in the presence of non-watertightness and other geometric errors.
@@ Appendix A Analytic discontinuity fix for Near-Field GWN
In this section, we provide a formal proof of the correction term in Equation 7. Assuming the surface and query are shifted such that the query is at the origin, one can compute the GWN in the “near-field” using the Stokes’ theorem boundary formulation, alongside a correction term which adds or subtracts $0.5$ for each intersection between the surface and the $z$-axis.
@@
For notational simplicity, we continue to refer to this deformed surface as $S$, such that $D\subset S$. To evaluate $w_{S}$, we now separate consider the relevant surface integral to be split between the disk $D$ and the remaining surface $S\setminus D$, such that
@@
Along with our handling of coincident points in Section 3.3, this yields Equation 7, as desired. $\square$
@@ Appendix B Pseudocode for Additional Algorithms
ALGORITHM 2 LinePatchIntersections Find all intersections between a line and a trimmed Bézier or NURBS surface.
@@
ALGORITHM 3 LineBezierIntersectionsRecursive Recursively find all intersections between a line and a Bézier surface. For nearly bilinear surfaces, we use the GARP algorithm (Reshetov, 2019) to directly find intersections.

### v3

@@ 4.1. Numerical Integration of Boundary Formulation
Comparison of numerical integration schemes for the 3D GWN. Comparisons between integration with a boundary formulation vs. tensor product rule, and fixed-order vs adaptive rules, valuated on slices of the top portion of a sphere. The plots for the fixed-order rules shows different spatial distributions of error. The plots for the adaptive rules show how both result in the same error, but with higher cost for the surface rule.
@@
We show the results of this comparison in the “Non-Adaptive Quadrature” row of Figure 9, where errors in the 1D scheme are only present near the surface boundary, while errors in the 2D scheme are present for all points near the surface. The accuracy of this 2D scheme, and in some sense even its applicability, would be further compromised by the presence of trimming curves, which are far more naturally handled in our boundary scheme.
The error in both methods can be mitigated through the use of adaptive quadrature, describing a “Boundary-adaptive” and “Surface-adaptive” scheme respectively. The performance of each method is shown in the “Adaptive Quadrature” row of Figure 9, showing that the surface-adaptive scheme would require thousands or even tens of thousands of times as many surface evaluations to achieve the same level of accuracy as our boundary-adaptive strategy, a cost which is further exacerbated when subdividing trimmed surfaces.
In (Martens and Bessmeltsev, 2025), this surface-adaptive scheme is referred to simply as “adaptive quadrature”, which largely explains their characterization that quadrature is inapplicable for complex CAD surfaces. In contrast, as can be seen in Figure 10, a boundary-adaptive strategy outperforms the reported metrics of that work on similar shapes, and is capable of doing so even without taking advantage of per-query parallelism, the planar characteristics of the input boundary curves, or any structure in the input set of query points. Figure 10 also demonstrates on practical examples how algorithm performance is influenced by the position of the query point. As expected, “near” points close to surface interiors have negligible overhead compared to “far” points, while only those close to a surface boundary are more expensive since they must often be treated as edge-cases.
@@
GWN fields and per-query evaluation times for a simple and complex shape The GWN field and per-query point evaluation times are shown for a simple shape and one with many more surfaces. In the plot of evaluation times, the most expensive points are near the surface boundary. The shape with more surfaces has increased evaluation time across the domain.
@@ 4.2. Accuracy Evaluation
We now consider the full implementation of our algorithm, which solves Equation 6 at arbitrary points up to arbitrary accuracy with a more sophisticated, error-controlled adaptive quadrature scheme. In Figure 11, we demonstrate the versatility of our algorithm by evaluating the GWN on several models adapted from the ABC dataset (Koch et al., 2019), a collection of CAD models of varying quality used to benchmark algorithms in computer graphics and machine learning. The specific models used in this work can be found in (Axom_data, 2025). In order to exaggerate the effects of the GWN field on open 3D shapes, we have intentionally made each of these shapes non-watertight through the deletion of several surfaces. We present qualitative results on additional watertight and non-watertight CAD models in the supplemental materials to this manuscript.
In general, a GWN-derived containment query is most useful in contexts where the objective is to faithfully represent input geometry, as opposed to when there is an unknown, ground-truth shape for which the provided boundaries are a subset. The GWN field itself is uniquely determined by the set of input surfaces, and so it cannot correct the geometric errors in a surface or its trimming curves, which are themselves assumed to be inextricably defined within the input. However, the subsequent containment query remains robust to such errors, as the containment result is not meaningfully changed outside a small region around the error.
This is demonstrated in Figure 11, where, for each example, the GWN field neatly partitions the domain into exterior and interior regions far from open boundaries, while providing reasonable containment decisions elsewhere. We can see this directly in the right-most column of Figure 11, which shows how rounding the scalar GWN field accounts for small errors in the surface by defining an implicit boundary through the $0.5$-isosurface.
@@
GWN evaluation for proposed method vs. triangulation vs. point cloud sampling GWN evaluation for proposed method vs. triangulation vs. point cloud sampling, seen on three increasingly zoomed in views of a complex, watertight CAD shape.
@@
In Figure 12 we compare our proposed algorithm to the above adaptations of conventional alternatives on subsets of a watertight “Gear” shape, with each view considered on a structured grid of $\times500$ sample points. In Figure 12 (top), we consider the shape when it is coarsely, yet still meaningfully, discretized according to each method to see the primary characteristics of the resulting GWN field. For the triangulation method, we see, as expected, that curved surfaces are made linear, resulting in unpredictable errors near the surface boundary while still producing integer-valued winding numbers for this watertight model.
In contrast, while the 0.5-isosurface of the scalar field provided by the point cloud method generally captures the curved features, the GWN field in its totality diverges significantly from the integer-valued ground truth. Although the dipoles correctly place nearby query points on the correct side of the curved surface, their large magnitude results in scalar GWN values outside [0.5,1.5][-0.5,1.5]. This would lead to further misclassifications if containment were determined with the conventional strategy of rounding, followed by an even-odd or non-zero rule, which might be necessary on more complex shapes with nested surfaces.
@@
To see this error more quantitatively, we record in Figure 12 (bottom) the total number of misclassifications on increasingly narrow views of the shape using the triangulation approach (left) and the point cloud approach (right), with each misclassified point in each $\times500$ grid marked in red. While the number of misclassifications for each discretization strategy (generally) decreases as the resolution of the discretization increases, there is no level at which no misclassifications occur. Further, although the linear deflection parameter controls how far the triangulation can differ from the original shape, in effect ensuring that no points within a fixed absolute distance of the surface will be misclassified, it can be difficult to know a priori what level of discretization is necessary for correct classification at user-specified query points. This is well-illustrated in the narrowest view, in which the increased deflection parameter from $10^{-3}$ m to $10^{-5}$ m does not necessarily increase the accuracy of the method, as all points which are misclassified are by coincidence still closer than $10^{-5}$ m to the surface. This relationship is further obfuscated in the case of a point cloud representation, where it is similarly difficult to gauge the appropriate sampling density for a point cloud that will lead to accurate classifications. In either case, this can lead to unnecessary computational expense in the form of globally dense triangulations or sampling procedures. In contrast, operating directly on the trimmed surface allows our method to perform adaptively for each query to the requested level of accuracy and correctly classify the containment of all query points.
@@
Shape Bobbin Trailer Pipe Spring Spring, Two-Patch
@@ 4.3. Performance Evaluation
Other differences come from more general characteristics of the component NURBS patches, which are best illustrated through comparison of the “Trailer” and “Pipe” examples. Although the Trailer shape has more surfaces, the NURBS patches themselves are predominantly flat with simple trimming curves. In contrast, the Pipe shape is more complicated, with many of its component NURBS patches being large spheres and cylinders which are trimmed and filleted. Thus, while the times to process the individual far- and near-field cases are comparable, resolving the edge-cases of the Pipe example is more challenging, and therefore more expensive. We present similar timing tests on other watertight CAD shapes from the ABC dataset in the supplemental materials to this manuscript.
@@ 5.1.1. Line-Surface Intersections
We use a geometric approach to identify the isoparametric coordinates of all signed intersections between a line and an untrimmed NURBS patch in the parameter space of the patch. This involves recursively subdividing the patch geometry (either through Bézier extraction of NURBS patches or bisection of individual Bézier patches) and clipping away portions of the surface where intersections provably cannot occur. By maintaining a record of each subdivision patch’s placement within the parameter space of the original surface, this method converges linearly to the set of parametric points $(u,v)$ at which the line intersects the untrimmed NURBS patch.
@@
Sensitivity study for line-surface intersection tolerance Example shape and line plot showing the sensitivity of accuracy to the tolerance for the line-surface intersection routine. As the tolerance decreases logarithmically, the number of randomly sampled points evaluated before any are misclassified increases to over ten million. The average evaluation cost increases slowly.
@@ 5.1.2. Adaptive Quadrature and Disk Radius
Sensitivity study for adaptive quadrature tolerance Three line plots showing the sensitivity of accuracy to the tolerance for the adaptive quadrature routine. As the tolerance decreases logarithmically, the number of randomly sampled points evaluated before any are misclassified increases to over ten million, while the cost of the procedure increases. As the disk radius increases from 0.1% to 1% to 10%, the cost increases, but the tolerance at which no misclassifications are observed decreases.
In our current implementation, memoized quadrature nodes cannot be reused if a patch’s trimming curves change during evaluation of the GWN. As a result, there is a strong correlation between overall computational performance and the choice of the disk radius used to identify and resolve near-field and edge-case configurations. On the one hand, a smaller disk radius reduces the number of edge-cases (thereby improving reuse of cached quadrature nodes), since the disk is less likely to intersect existing trimming curves. On the other hand, a smaller disk radius allows query points to be closer to the singularity in the boundary curve integrals, necessitating a stricter tolerance for $\epsilon_{q}$ to ensure consistent accuracy.
To demonstrate how both the disk radius (measured as a percent of each patch’s parameter space bounding box diagonal) and $\epsilon_{q}$ affects the accuracy and computational efficiency of the algorithm, we repeat the numerical experiment for $\epsilon_{ls}$ in Section 5.1.1 and report the results in Figure 14, varying $\epsilon_{q}$ during the evaluation of the GWN for disks of various sizes. In these examples, we set $\epsilon_{ls}=10^{-6}$, and further note that $\epsilon_{ls}$ and the disk radius have an independent effect on the method’s accuracy. We see that as the radius decreases, a more strict tolerance for $\epsilon_{q}$ is needed to ensure there are exceptionally few misclassifications. Furthermore, we see that across disk radii, a stricter tolerance involves a greater computational cost, but this cost is marginal relative to the overall effect of disk radius. It is for this reason that we use for our algorithm a disk radius that is 1% of the parameter space bounding box diagonal, as it is small enough to allow full utilization of the memoization scheme, while still allowing for a somewhat lenient tolerance of $\epsilon_{q}=10^{-6}$.
@@ 5.1.3. Disk Extraction
Disk extraction via trimming curves in parameter space.
@@ 5.2. Robustness to Defective Trimming Curves
Demonstration of algorithm robustness to invalid trimming curves The GWN field and containment classifications for four instances of trimming curve geometry manipulated to be invalid. Trimming curves are shown to be perturbed and non-watertight at varying levels of severity. With increasing severity the GWN field becomes more non-smooth, but the containment decisions are virtually unaffected.
@@
First is robustness to errors between the (valid) trimmed NURBS surfaces in the model, which is necessary when numerical gaps and overlaps between adjacent patches make conventional containment techniques like ray casting inaccurate. By the linearity of the GWN, our algorithm is evaluated independently for each patch, ensuring our containment decision is reasonable for this kind of messy input. We achieve this form of robustness through accurate and precise calculation of the GWN field at arbitrary 3D points, one surface at a time. This property of our algorithm is exhibited on watertight and non-watertight examples throughout this paper (e.g. Figures 11011 and 12) as well as in the supplement.
@@
Figure 16 demonstrates how these decisions impact our calculation of the GWN field in the presence of a surface with poorly defined trimming curves, and in turn how this GWN field impacts the subsequent containment decision. To the original shape in Figure 16 (top left), we perturb the trimming curves of the front face while keeping adjacent curves connected. The resulting set of trimming curves, shown in Figure 16 (bottom left), is still closed and connected, but the perturbation has introduced self-intersections that form small clockwise loops. The boundary formulation itself is indifferent to trimming curve orientation, and so for query points away from these self-intersections, Algorithm 1 continues to produce a GWN that degrades smoothly around the mismatch with the new boundary. While this type of trimming curve geometry poses a challenge for our correction term (Equation 6) and edge-case strategy, resulting in a noisier GWN field, the containment decisions remain consistent with those of the original model.
Next, we consider collections of non-watertight trimming curves. For Figure 16 (top right) we delete trimming curves at each corner in parameter space, and for Figure 16 (bottom right) we delete a middle portion from each curve. In the presence of these defective trimming curves, we observe that the computed GWN field does not degrade smoothly in the same way that it does around open, yet well-defined edges. For example, applying our boundary formulation to the same malformed surface returns slightly different GWN values at the same query point depending on the chosen form of the line integral. This is seen most prominently in the boundary between regions handled by different cases, such as between far-field and edge-case scenarios. Nevertheless, we see that these discontinuities are concentrated near geometric errors, such that rounding and applying an even-odd rule to the calculated GWN field still achieves excellent containment decisions for these examples.
@@ 6. Conclusions
In this work, we have introduced the first algorithm to accurately and precisely compute generalized winding numbers over unstructured collections of trimmed NURBS patches. We have used this to define a robust containment query that can be applied to general CAD models, even in the presence of non-watertightness, overlapping boundary components, and other geometric errors.
@@ Appendix A Analytic discontinuity fix for Near-Field GWN
In this section, we provide a formal proof of the correction term in Equation 6. Assuming the surface and query are shifted such that the query is at the origin, one can compute the GWN in the “near-field” using the Stokes’ theorem boundary formulation, alongside a correction term which adds or subtracts $0.5$ for each signed intersection between the surface and the $z$-axis.
@@
For notational simplicity, we continue to refer to this deformed surface as $S$, such that $D\subset S$. To evaluate $w_{S}$, we now consider the relevant surface integral to be split between the disk $D$ and the remaining surface $S\setminus D$, such that
@@
Along with our handling of points coincident with the surface in Section 3.3, this yields Equation 6, as desired. $\square$
@@ Appendix B Pseudocode for Additional Algorithms
ALGORITHM 2 LinePatchIntersections Find all intersections between an oriented line and a trimmed Bézier or NURBS surface.
@@
ALGORITHM 3 LineBezierIntersectionsRecursive Recursively find all intersections between an oriented line and a Bézier surface. For nearly bilinear surfaces, we use the GARP algorithm (Reshetov, 2019) to directly find intersections.